# A quasi-one-dimensional hydrogen-bonded monolayer ice phase

### v1

@@ Appendix C Hydrogen bonding characterization of the hexagonal and pentagonal phases
Figure 8: a) The log-probability and number of hydrogen bonds in $(\theta,\phi)$-space for the hexagonal phase. b) The same plots for the pentagonal phase.
@@
In Fig. 2(b), we provide 2D log-probability profiles in $\phi$ and $\theta$ space for molecules in the flat-rhombic phase. We also provide the average number of hydrogen bonds observed in the flat-rhombic phase in $\phi$ and $\theta$ space. Surprisingly, the maxima in the log-probability profiles do not align with the maxima in the number of hydrogen bonds plot. This suggested that the number of hydrogen bonds alone does not predict the stability of a nanoconfined ice configuration. Fig. 8 shows the same types of plots for the hexagonal and pentagonal phases. These plots serve as a baseline for our expectations from bulk ice, where the Bernal-Fowler ice rules would dictate that bulk ice configurations that satisfy the maximum hydrogen bonding coordination of 4 would be the most stable ordered configurations. The plots in Fig. 8 for the hexagonal and pentagonal phases exhibit exact alignment between the log-probability maxima and the number of hydrogen bond maxima, which suggests that this hydrogen bonding-oriented picture holds for these phases.
Appendix D Dielectric nature of the flat-rhombic phase.
We analyzed the in-plane $\varepsilon_{\parallel}$ and out-of-plane $\varepsilon_{\perp}$ dielectric constants of nanoconfined water based on the variance of a linear polarization model: the TIP4P point charge model [74]. We observe a net dipole moment of the water molecules along the direction of the quasi-one-dimensional chains and low and near-constant in-plane and out-of-plane dielectric constants, in agreement with the experiments in Ref. 50. The singularity observed in Fig. 9 corresponds to the transition to the hexatic phase [16], which exhibits an increased in-plane dielectric constant but a similarly small out-of-plane dielectric constant.
Figure 9: The dielectric constant of the flat-rhombic phase across the range of temperatures considered in this work. The singularity at 380 K corresponds to the transition to the hexatic phase of nanoconfined water [55].
Appendix E Impact of quantum nuclear effects on the flat-rhombic phase
@@
Fig. 10(b) shows the log-probability and average number of hydrogen bonds in $(\theta,\phi)$-space for the PIMD simulations. Fig 10(a) just serves as a convenient replication of the results in Fig. 2(b) for comparison. The major result, i.e., the log-probability maxima not aligning with the number of hydrogen bond maxima, holds in the presence of quantum nuclear motion, suggesting that hydrogen bonds are not the primary stabilization interaction in the flat-rhombic phase. However, the PIMD simulations exhibit a broader exploration of $(\theta,\phi)$-space at all temperatures. In particular, they enhance the in-plane rotational motion of the molecules about its dipole vector.
@@
Figure 10: a) The classical log-probability profiles and number of hydrogen bond plots in $(\theta,\phi)$-space. These are the same as in Fig. 2(b). b) The same plots but computed for PIMD simulation trajectories that account for quantum nuclear motion.
@@
We also compute the free energy profile along $\sigma$ in these PIMD simulations in Fig. 11. Comparing Fig. 11 to Fig. 4(c) shows that the classical and PIMD simulations exhibit the same qualitative behaviour. However, the free energy barriers for the $\sigma$-switching motion are lower due to zero-point fluctuations.
@@
Figure 11: Free energy profiles for the $\sigma$ order parameter computed from PIMD simulations at various temperatures. This can be directly compared to Fig. 4(c) in the main text.

### v2

@@ V Hydrogen Bonding Characterization of the Hexagonal and Pentagonal Phases
Fig. 12: |||| Molecular orientations in the hexagonal and pentagonal phases. a) The log-probability and number of hydrogen bonds in $(\theta,\phi)$-space for the hexagonal phase. b) The same plots for the pentagonal phase.
@@
In Fig. 3 of the main text, we provide 2D log-probability profiles in $\phi$ and $\theta$ space for molecules in the flat-rhombic phase. We also provide the average number of hydrogen bonds observed in the flat-rhombic phase in $\phi$ and $\theta$ space. Surprisingly, the maxima in the log-probability profiles do not align with the maximum in the number of hydrogen bonds plot. This suggested that the number of hydrogen bonds alone does not predict the stability of a nanoconfined ice configuration. Supplementary Figure 12 shows the same types of plots for the hexagonal and pentagonal phases. These plots serve as a baseline for our expectations from bulk ice, where the Bernal-Fowler ice rules would dictate that bulk ice configurations that satisfy the maximum hydrogen bonding coordination of 4 would be the most stable ordered configurations. The plots in Supplementary Figure 12 for the hexagonal and pentagonal phases exhibit exact alignment between the log-probability maxima and the number of hydrogen bond maxima, which suggests that this hydrogen bonding-oriented picture holds for these phases.
VI Impact of Quantum Nuclear Effects on the Flat-Rhombic Phase
@@
Supplementary Figure 13(b) shows the log-probability and average number of hydrogen bonds in $(\theta,\phi)$-space for the PIMD simulations. Supplementary Figure 13(a) just serves as a convenient replication of the results from Fig. 3(b) of the main text for comparison. The major result, i.e., the fact that the log-probability maxima do not align with the maximum in the number of hydrogen bond plot, holds in the presence of quantum nuclear motion, suggesting that hydrogen bonds are not the primary stabilization interaction in the flat-rhombic phase. However, the PIMD simulations exhibit a broader exploration of $(\theta,\phi)$-space at all temperatures. In particular, they enhance the in-plane rotational motion of the molecules about its dipole vector.
@@
Fig. 13: |||| Nuclear quantum effects and molecular orientations. a) The classical log-probability profiles and number of hydrogen bond plots in $(\theta,\phi)$-space. b) The same plots but computed for PIMD simulation trajectories that account for quantum nuclear motion.
@@
We also compute the free energy profile along $\sigma$ in these PIMD simulations in Supplementary Figure 14. Comparing Supplementary Figure 14 to Fig. 6(c) of the main text shows that the classical and PIMD simulations exhibit the same qualitative behavior. However, the free energy barriers for the $\sigma$-switching motion are lower due to zero-point fluctuations, and the minima in the free energy profiles move closer to $\sigma=0$ due to the increased proton disorder, as discussed in the main text.
@@
Fig. 14: |||| Nuclear quantum effects and the $\sigma$ order parameter. Free energy profiles for the $\sigma$ order parameter computed from PIMD simulations at various temperatures. This can be directly compared to Fig. 6(c) in the main text.